# LCI: a Lightweight Communication Interface for Efficient Asynchronous Multithreaded Communication

### v1

@@ 3.2.5. Operation Return Values
  • retry: The operation needs to be resubmitted due to temporary resource unavailability.
  • fatal error: The operation has failed due to a fatal error.
Fatal errors are reported through C++ exceptions. The returned status_t object reports the other three categories. Each category includes multiple error codes to deliver more information (e.g., what resource is temporarily unavailable). When the status is done, the returned status object contains valid information about the completed operation.
Line 39-44 shows how iRPCLib handles these return values. It just returns false if it gets a retry error (Line 39). In this case, the upper layer can do something meaningful, such as polling other task queues or aggregating RPC messages. If the communication is immediately completed, the return status object will contain valid information, and iRPCLib just manually invokes send_cb (Line 41).
Compared to the binary return values of MPI nonblocking operations, the additional done and retry provide more informative feedback, enabling finer-grained control and potentially unlocking further optimization opportunities.
Completion Checking
@@
Under the hood, a completion object is a functor with a virtual signal method that takes a status_t object as an argument. Derived from it, LCI defines four built-in completion object types: handler, queue, synchronizer, and graph. Synchronizer is similar to MPI requests but can accept multiple signals before becoming ready. Graph is a more advanced completion object type similar to CUDA Graph (NVIDIA, 2019) that allows users to specify a set of communication operations or user-provided functions with a partial execution order. If operation $u$ precedes operation $v$ in that order, then $v$ will be started only after $u$ completes. The local partial execution order and the ordering imposed by communication operations allow intuitive implementations of complex nonblocking collective algorithms.
@@ 3.2.6. Progress
In MPI, communication progressing happens as a side effect of certain MPI calls (typically MPI_Test* and all blocking functions). In contrast, LCI defines an explicit progress function. Users can select whether progress is invoked by a distinct thread or as a side-effect of other operations and how frequently progress should be called. Line 70 shows how the backend layer uses the OFF version of the progress function to make progress on the thread-local device.
@@ 3.3.1. Other Advanced Features
Listing 2 assumes the upper layer supplies plain send buffers, and LCI also uses plain buffers to deliver incoming active messages. Alternatively, advanced users can explicitly ask LCI for packets and directly assemble the message in it. They can also instruct LCI to deliver incoming active messages in packets. These practices can save memory copy for buffer-copy protocol.
In addition, LCI follows the common practice of many low-level communication libraries by providing an explicit memory registration function. Memory registration is optional for local buffers but mandatory for remote buffers. LCI supports on-demand paging when the underlying hardware allows it.
Besides a single source and target buffer, LCI also supports transmitting a list of source and target buffers in a single communication posting operation to reduce the overheads related to request posting, handshakes, and completion signaling.
@@ 3.3.2. Send-Receive Semantics
LCI adopts the send-receive semantics proposed in (Dang et al., 2016), namely, out-of-order delivery and restricted wildcard matching, to avoid sequential bottlenecks inside the runtime. The in-order delivery and wildcard matching have long been seen as a stumbling block for efficient multithreaded MPI implementation, as they require centralized matching queues that are hard to parallelize. Weakening them allows LCI to adopt a more efficient hashtable-based matching engine. By default, LCI matches send and receive by the (matching engine, source rank, tag) tuple on the target side. Users can still achieve in-order matching for send-receives by encoding ordering information into the tag field. They can also set the matching_policy optional argument to tag_only or rank_only when posting sends and receives to achieve wildcards similar to those of MPI, except that the sender needs to know the sent message will be matched by a wildcard receive call. Under the hood, the matching_policy will instruct the matching engine on how to make the insertion key based on rank and tag. Users can also achieve more flexible matching policies by supplying their own make_key function.
@@ 4. LCI Runtime
Communication activities inside the LCI runtime are carefully decomposed into operations of multiple independent resources, while each resource is carefully optimized with threading efficiency in mind. Key optimizations include atomic-based data structures, fine-grained locking, thread-local storage, and try lock wrappers.
@@ 4.1.1. Prerequisite: Multi-Producer-Multi-Consumer (MPMC) Array
We find it a common need for LCI to store certain resources in an array for future reference. Such arrays are rarely written but frequently read, and the array size is usually unknown at compilation time. For example, the registered completion object array is only written (appended) during a new registration (usually not on the critical path) but is read whenever an active message or RMA with notification message is received. We do not want to preallocate a large array as it may waste memory and restrict the total number of registered completion objects.
To meet this need, we implement a simple MPMC array that supports dynamic resizing and fast read. It borrows the key idea in (Alexandrescu and Michael, 2004): a write and append (and the potential resize) is protected by a lock to prevent missed writes, but read is lock-free. Every resize swaps the old array with a new one that doubles the size, and the deallocation of the old array is postponed to prevent the read from reading invalid memory.
@@ 4.1.2. Packet Pool
The packet pool is responsible for efficient allocation (get) and deallocation (put) of fixed-sized pre-registered buffers, which we call packets. get can be nonblocking and will return a nullptr when it fails the first packet stealing attempts (and post_comm returns retry). The packet pool is implemented as a collection of thread-local double-ended queues (deque). The list of thread-local deques is managed by an MPMC array. By default, every thread puts and gets packets from its own deque. When the local deque is empty, the thread will try stealing half of the total packets from a randomly selected deque. Local packet put and get are performed at the tail end, and packet stealing is performed at the head end to achieve better cache locality. Thread safety is achieved with a per-deque spinlock, so there should be no thread contention during normal operation.
@@ 4.1.3. Matching Engine
The matching engine is responsible for matching the incoming sends with user-posted receives at the target side. It contains two major methods: make_key generates a matching key based on source rank, tag, and user-supplied matching_policy; insert tries inserting a key-value pair with a type (send or receive) and will either return 0, meaning the entry has been inserted, or the matched values if an entry with the same key and a complementary type has been found. The default implementation is based on a hashtable where each bucket is a list of queues. Thread safety is achieved with a per-bucket spinlock, and we do not expect severe thread contention, given that the bucket number (by default 65536) is significantly larger than the thread number (on the order of tens to hundreds). Special optimization is applied to the case when a bucket contains no more than three queues and when a queue contains no more than two sends or receives, where we use fixed-size arrays instead of linked lists for the buckets/queues. Therefore, when the load factor is low, the hashtable can perform an insertion with a single cache miss.
@@ 4.1.4. Completion Objects
All LCI built-in completion objects are atomic-based. Synchronizer is implemented as an atomic flag (when expecting one signal) or a fixed-sized array protected by two atomic counters (when expecting multiple signals). Completion queue has two implementations: one based on the state-of-the-art LCRQ (Morrison and Afek, 2013) and the other based on a hand-written Fetch-And-Add-based fix-sized array. Completion handler is essentially a function and does not need any special treatment. Every node in the completion graph uses an atomic counter to track the number of received signals. Every ready node will be immediately fired, and a completed node will signal all its descendants.
@@ 4.1.5. Backlog Queue
The backlog queue is used to store communication requests that cannot be immediately submitted and cannot be back-propagated to the user. For example, when the progress engine wants to send a handshake message, but the underlying network send queue is full. LCI expects such scenarios to be rare, so we implement it with a simple C++ queue with a spinlock. An atomic flag prevents the progress engine from unnecessarily polling an empty backlog queue.
@@ 4.2.1. The Network Backend Layer
LCI isolates different network backends from its core runtime with a simple network backend wrapper. The backend abstraction operates on two resources: network context and network device. Each LCI runtime maps to a network context, which contains global network resources. Each LCI device maps to a network device, which contains network resources accessed on the critical path.
All communication operations on the critical path are posted to a network device. These operations include posting network-layer send/recv/write/read, polling for completed operations, and (de)registering memory. LCI does not require the ability to handle tag matching and unexpected receive from the network backends. The LCI progress engine ensures there are always enough pre-posted receives in the device. LCI expects two threads operating on different network devices not to interfere with each other.
@@
4.2.2. Trylock Wrapper
Lower-level network stacks such as ibv and ofi generally use spin locks to ensure thread safety and usually blockingly acquire them. To mitigate the cost of blocking on these locks, we examine the backend source code to identify the lock granularity and wrap all corresponding accesses with a try lock. For example, an ibv completion queue is protected by a spin lock, so we create a spin lock for each ibv completion queue at LCI layer and try_lock the corresponding LCI-layer lock before we access the ibv completion queue through (ibv_poll_cq). If the try lock fails, we will return the retry error code to the caller. This gives LCI clients more optimization opportunities during network contention.
@@ 4.2.2. Trylock Wrapper
libibverbs is the lowest-level public API for Infiniband. It can also be run on top of High-speed Ethernet devices through RDMA over Converged Ethernet (RoCE). We focus on its mlx5 provider here as it is the latest and most widely used one. Each libibverbs queue pair, shared receive queue, and completion queues are protected by their own spinlock. In addition, each queue pair is associated with a set of hardware resources (micro User Access Region or uUAR) that are protected by its own lock on the host side. Different queue pairs may share the same uUAR (Zambre et al., 2018). libibverbs users can use thread domains to explicitly associate queue pairs with uUARs. The memory (de)registration functions do not acquire any locks.
The LCI ibv backend puts an ibv completion queue, an ibv shared receive queue, and a collection of ibv queue pairs in a network device. LCI uses a try lock wrapper for every ibv completion queue and shared receive queue. An LCI device attribute ibv_td_strategy controls the way LCI uses thread domains. By default, it will create a thread domain for every ibv queue pair (the per_qp strategy). Users can also ask LCI to allocate a single thread domain for all queue pairs of a device (the all_qp strategy) or not use thread domains at all (the none strategy). The all_qp strategy is recommended when each thread has a dedicated LCI device. LCI uses a try lock wrapper for every queue pair in the per_qp case and uses a try lock wrapper for all queue pairs of the device in the other two cases.
With libibverbs, LCI can provide a contention-free guarantee not only for threads operating on different devices but also for threads operating on different ibv data structures (queue pairs, completion queues, shared receive queues). This means there will be no interference between a worker thread posting communication and a background thread progressing the communication, which is typical in asynchronous programming systems such as AMTs (Bauer et al., 2 11).
@@
libfabric is a portable low-level network API that supports many network providers. It is also currently the lowest-level public API for HPE Slingshot-11. The LCI ofi backend is designed with the libfabric cxi provider and verbs provider in mind. Both providers have similar lock granularity: every endpoint has a single spin lock; all post_send/recv on the endpoint and poll_cq on associated completion queues need to acquire the endpoint lock; the memory (de)registration function involves the use of a registration cache, which is allocated per domain and is protected with a pthread mutex.
The LCI ofi backend, therefore, puts in a network device an ofi domain, endpoint, and completion queue. It uses a single try lock wrapper for the endpoint. It does not employ a try lock wrapper for the memory (de)registration functions, as LCI does not yet have a way to backpropagate the memory registration failure to the user.
In general, the current libfabric provider implementation has a coarser lock granularity than libibverbs, which makes it less efficient in multi-threading scenarios. However, the libfabric interface is general enough to accommodate additional optimizations in the provider implementation. libfabric defines a more advanced FI_THREAD_FID threading support level that only requires serialization to individual libfabric objects. Combined with libfabric’s scalable endpoint, it could achieve lock granularity similar to that of libibverbs. However, current providers do not exploit this threading support level with additional optimizations. The two providers also do not support the scalable endpoint feature.
@@ 4.3. Communication Protocol
LCI adopts communication protocols similar to existing communication libraries, so we will briefly mention them due to page limit. For the send-receive and active message operations, depending on the message size, LCI adopts three different communication protocols: inject, buffer-copy, and zero-copy. For put/get operations, LCI directly translates them into the corresponding low-level network operations. Due to the lack of support for RDMA read with notification in the interconnects we have access to, LCI does not implement the get with signal communication operation for the time being.
@@ 4.4. Putting Everything Together
Figure 1 shows an overview of the LCI runtime architecture. When the user posts a communication, (1) if it is a receive, a receive descriptor will be inserted into the matching engine; (2) otherwise, the communication request will be posted to the device. When the user invokes the progress engine, it will (3) first check the backlog queue and retry the communication requests in that queue; and (4) poll the device for completed operation and react accordingly. The reaction may involve (5) inserting an incoming send into the matching table, (6) signaling a completion object, (7) replenishing the pre-posted receives, or (8) posting another communication request to the device as part of the rendezvous (zero-copy) protocol. When either the communication posting procedure or the progress engine finds a match in the matching engine, it will either (9) signal the completion object or (10) post another communication to continue the rendezvous protocol. (11) The completion checking procedure will query the completion object for the status of posted communication.

### v2

@@ 4.2.5. Operation Return Values
  • retry: The operation must be resubmitted due to temporary resource unavailability.
  • error: The operation has failed due to an error.
Errors are reported through C++ exceptions. Applications can catch these exceptions and continue execution. The returned status_t object reports the other three categories. Each category includes multiple status codes to deliver more information (e.g., what resource is temporarily unavailable). When the status is done, the returned status object contains valid information about the completed operation.
Line 39-44 shows how iRPCLib handles these return values. It just returns false if it gets a retry error (Line 39). In this case, the upper layer can do something meaningful, such as polling other task queues or aggregating these messages. If the communication is immediately completed, the returned status object will contain valid information, and iRPCLib just manually invokes send_cb (Line 41).
Compared to the binary return values of MPI nonblocking operations, the additional done and retry eliminate the blocking retrying loop inside common MPI_Isend implementations and unlock more optimization opportunities for applications.
4.2.6. Completion Checking
@@
Under the hood, a completion object is a functor with a virtual signal method that takes a status_t object as an argument. Derived from it, LCI defines five built-in completion object types: handler, queue, counter, synchronizer, and graph. Counter simply records the number of times it has been signaled. Synchronizer is similar to MPI requests but can accept multiple signals before becoming ready. Graph is a more advanced completion object type conceptually similar to CUDA Graph (NVIDIA, 2019) that allows users to specify a set of communication operations or user-provided functions with a partial execution order. If operation $u$ precedes operation $v$ in that order, then $v$ will be started only after $u$ completes. The local partial execution order and the ordering imposed by communication operations allow intuitive implementations of complex nonblocking collective algorithms. Users can also define their own custom completion types.
@@ 4.2.7. Progress
In MPI, communication progress happens as a side effect of certain MPI calls selected by the implementer (typically MPI_Test* and all blocking functions). In contrast, LCI offers an explicit progress function. Users can decide whether progress is invoked by a distinct thread or as a side effect of other operations and how frequently progress should be called. Line 69 shows how the backend layer uses the named-parameter version of the progress function to make progress on the thread-local device.
@@ 4.3.1. Other Advanced Features
Listing 2 assumes the upper layer supplies plain send buffers, and LCI also uses plain buffers to deliver incoming active messages. Alternatively, advanced users can explicitly ask LCI for packets and directly assemble the message in them. They can also instruct LCI to deliver incoming active messages in packets. These practices can save memory copy for the buffer-copy protocol.
In addition, LCI follows the common practice of many low-level communication libraries by providing an explicit memory registration function. Memory registration is optional for local buffers but mandatory for remote buffers. LCI can also use on-demand paging when the underlying hardware supports it.
@@ 4.3.2. Send-Receive Semantics
LCI adopts the send-receive semantics proposed in (Dang et al., 2016), namely, out-of-order delivery and restricted wildcard matching, to avoid sequential bottlenecks inside the runtime. The in-order delivery and wildcard matching have long been seen as a stumbling block for efficient MPI_THREAD_MULTIPLE implementation, as they require centralized matching queues that are hard to parallelize. Relaxing them allows LCI to adopt a more efficient hashtable-based matching engine. By default, LCI matches send and receive by the (matching engine, source rank, tag) tuple on the target side. Users can still achieve in-order matching for send-receives by encoding ordering information into the tag field. LCI supports MPI-like wildcard matching, but the sender must also inform the runtime that the message will be matched by a wildcard receive. This is done by setting the optional argument matching_policy to tag_only or rank_only in both post_send and post_recv. Under the hood, the matching_policy will instruct the matching engine on how to make the insertion key based on rank and tag. Users can also achieve more flexible matching policies by supplying their own make_key function.
@@ 5. LCI Runtime
Communication activities inside the LCI runtime are carefully decomposed into operations of multiple independent components, while each component is carefully optimized with threading efficiency in mind. Key optimizations include atomic-based data structures, fine-grained locking, thread-local storage, and try-lock wrappers. LCI refers to these components as resources.
@@ 5.1.1. Prerequisite: Multi-Producer-Multi-Consumer (MPMC) Array
We find it a common need for LCI to store certain resources in an array for future reference. Such arrays are rarely written but frequently read, and the array size is usually unknown at compilation time. For example, the remote completion handle array is only written (appended) during a new completion object registration (usually not on the critical path), but is read whenever an active message or RMA with notification message is received. We do not want to preallocate a large array as it may waste memory and restrict the total number of remote completion handles.
We implement a simple MPMC array that supports dynamic resizing and fast read to meet this need. It borrows a key idea from (Alexandrescu and Michael, 2004): a write and append (and the potential resize) is protected by a lock to prevent missed writes, but read is lock-free. Every resize swaps the old array with a new one that doubles the size, and the deallocation of the old array is postponed to prevent the read from reading invalid memory.
@@ 5.1.2. Packet Pool
The packet pool is responsible for efficient allocation (get) and deallocation (put) of fixed-sized pre-registered buffers, referred to by LCI as packets. get can be nonblocking and will return a nullptr when it fails the first packet stealing attempts (and post_comm returns retry). The packet pool is implemented as a collection of thread-local double-ended queues (deque). An MPMC array manages the list of thread-local deques. By default, every thread puts and gets packets from its own deque. When the local deque is empty, the thread will try stealing half of its packets from a randomly selected deque. Local packet put and get are performed at the tail end, and packet stealing is performed at the head end to achieve better cache locality. Thread safety is achieved with a per-deque spinlock, so there should be no thread contention during normal put and get.
@@ 5.1.3. Matching Engine
The matching engine is responsible for matching the incoming sends with user-posted receives at the target side. It contains two major methods: make_key generates a matching key based on source rank, tag, and user-supplied matching_policy; insert tries inserting a key-value pair with a type (send or receive) and will either return 0, meaning the entry has been inserted, or the matched values if an entry with the same key and a complementary type has been found. The default implementation is based on a hashtable where each bucket is a list of queues. Thread safety is achieved with a per-bucket spinlock, and we do not expect severe thread contention, given that the bucket number (by default 65536) is significantly larger than the thread number (on the order of tens to hundreds). Special optimization is applied when a bucket contains no more than three queues and a queue contains no more than two sends or receives, where we use fixed-size arrays instead of linked lists for the buckets and queues. Therefore, when the load factor is low, the hashtable can perform an insertion with a single cache miss.
@@ 5.1.4. Completion Objects
All LCI built-in completion objects are atomic-based. Counter is implemented as an atomic integer. Synchronizer is implemented as an atomic flag (when expecting one signal) or a fixed-sized array protected by two atomic counters (when expecting multiple signals). Completion queue has two implementations: one based on the state-of-the-art LCRQ (Morrison and Afek, 2013) and the other based on a hand-written Fetch-And-Add-based fixed sized array. Completion handler is essentially a function pointer and does not need any special treatment. Every node in the completion graph uses an atomic counter to track the number of received signals. Every ready node will be immediately fired, and a completed node will signal all its descendants.
@@ 5.1.5. Backlog Queue
The backlog queue is used to store communication requests that cannot be immediately submitted and cannot be back-propagated to the user. For example, when the progress engine wants to post a network send (e.g., in a rendezvous protocol), but the underlying network send queue is full. Keep retrying such sends inside the progress engine may cause deadlocks, so these operations are pushed to the backlog queue instead. LCI expects such scenarios to be rare, so we implement it with a simple C++ queue with a spinlock. An atomic flag prevents the progress engine from unnecessarily polling an empty backlog queue.
@@ 5.2.1. The Network Backend Layer
LCI isolates different network backends from its core runtime with a simple network backend wrapper. The backend abstraction operates on two resources: network context and network device. Each LCI runtime maps to a network context, while each LCI device maps to a network device. A network device contains network resources accessed on the critical path.
All communication operations on the critical path are posted to a network device. These operations include posting network-layer send/recv/write/read, polling for completed operations, and (de)registering memory. LCI does not require the ability to handle tag matching and unexpected receive from the network backends. The LCI progress engine always ensures there are enough pre-posted receives in the device. LCI expects two threads operating on different network devices not to interfere with each other.
@@
5.2.2. Try-lock Wrapper
Lower-level network stacks such as ibv and ofi generally use spin-locks to ensure thread safety, and they are usually blocked while acquiring them. To mitigate the cost of blocking on these locks, we examine the backend source code to identify the lock granularity and wrap all corresponding accesses with a try-lock. For example, an ibv completion queue is protected by a spin-lock, so we create a spin-lock for each ibv completion queue at the LCI layer and try_lock the corresponding LCI-layer lock before we access the ibv completion queue through ibv_poll_cq. If the try-lock fails, we will return the retry error code to the caller. This gives LCI clients more optimization opportunities during network contention.
@@ 5.2.2. Try-lock Wrapper
libibverbs is the lowest-level public API for Infiniband. It can also be run on top of high-speed Ethernet devices through RDMA over Converged Ethernet (RoCE). We focus on its mlx5 provider here as it is the latest and most widely used one. Each libibverbs queue pair, shared receive queue, and completion queue is protected by its own spinlock. In addition, each queue pair is associated with a set of hardware resources (micro User Access Region or uUAR) that are protected by their own lock on the host side. Different queue pairs may share the same uUAR (Zambre et al., 2018). libibverbs users can use thread domains to explicitly associate queue pairs with uUARs. The memory (de)registration functions do not acquire any locks in user space.
The LCI ibv backend puts an ibv completion queue, an ibv shared receive queue, and a collection of ibv queue pairs in a network device. LCI uses a try-lock wrapper for every ibv completion queue and shared receive queue. An LCI device attribute ibv_td_strategy controls how LCI uses thread domains. By default, it will create a thread domain for every ibv queue pair (the per_qp strategy). Users can also ask LCI to allocate a single thread domain for all queue pairs of a device (the all_qp strategy) or not use thread domains at all (the none strategy). The all_qp strategy is recommended when each thread has a dedicated LCI device. LCI uses a try-lock wrapper for every queue pair in the per_qp case and a try-lock wrapper for all queue pairs of the device in the other two cases.
With libibverbs, LCI avoids contention not only for threads operating on different devices but also for threads operating on different ibv data structures (queue pairs, completion queues, shared receive queues). This means there will be no interference between a worker thread posting communication and a background thread progressing the network, typical in asynchronous programming systems such as AMTs (Bauer et al., 2 11).
@@ 5.2.4. Libfabric Analysis
Libfabric is a portable low-level network API that supports many network providers. It is also currently the lowest-level public API for HPE Slingshot-11. The LCI ofi backend is designed with the libfabric cxi provider and verbs provider in mind. Both providers have similar lock granularity: every endpoint has a single spin-lock; all post_send/recv on the endpoint and poll_cq on associated completion queues need to acquire the endpoint lock; the memory (de)registration function involves the use of a registration cache, which is allocated per domain and is protected with a global pthread mutex. To make matters worse, there is no way to disable the usage of the registration cache in the libfabric interface, and the cxi provider appears to consult the registration cache for almost every communication operation.
The LCI ofi backend puts in a network device an ofi domain, endpoint, and completion queue. It uses a single try-lock wrapper for each device (except its memory (de)registration functions) to mitigate the overhead of the libfabric endpoint spin-lock. We have not found a way to mitigate the impact of the global mutex for the internal registration cache, and it has been a major performance bottleneck in many of our evaluations.
In general, the per-endpoint spinlock and the global registration cache mutex make libfabric less efficient in multi-threading scenarios. However, the libfabric interface is general enough to accommodate additional optimizations in the provider implementation. The global registration lock can be optimized into a private lock for each registration cache. Moreover, libfabric defines a more advanced FI_THREAD_FID threading support level that only requires serialization to individual libfabric objects. Combined with libfabric’s scalable endpoint, it could achieve lock granularity similar to libibverbs. Currently, providers do not exploit this threading support level with additional optimizations. The two providers also do not support the scalable endpoint feature.
@@ 5.3. Communication Protocol
LCI adopts communication protocols similar to existing communication libraries, so we will briefly mention them due to the page limit. For the send-receive and active message operations, depending on the message size, LCI adopts three different communication protocols: inject, buffer-copy, and zero-copy (rendezvous). For put/get operations, LCI directly translates them into the corresponding low-level network operations. Due to the lack of support for RDMA read with notification in the interconnects we have access to, LCI does not implement the get with signal communication operation for the time being.
@@ 5.4. Putting Everything Together
Figure 2 shows an overview of the LCI runtime architecture. When the user posts a communication, (1) if it is a receive, a receive descriptor will be inserted into the matching engine; (2) otherwise, the communication request will be posted to the device. When the user invokes the progress engine, it will (3) first check the backlog queue and retry the communication requests in that queue; and (4) poll the device for completed operations and react accordingly. The reaction may involve (5) inserting an incoming send into the matching engine, (6) signaling a completion object, (7) replenishing the pre-posted receives, or (8) posting another communication request to the device as part of the rendezvous protocol. When either the communication posting procedure or the progress engine finds a match in the matching engine, it will either (9) signal the completion object or (10) post another communication to continue the rendezvous protocol. (11) The completion checking procedure will query the completion object for the status of posted communication.